OpenStudy (anonymous):

Find the sum and product of the roots. 3x2 - 5x + 7= 0

OpenStudy (callisto):

For a quadratic equation \(ax^2 + bx +c =0 \), sum of roots = -b/a product of roots = c/a Identify the value of a, b and c in the given equation \(3x^2 -5x + 7 = 0 \) and plug in the value to get your answer.

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

welcome.